# End-to-End Framework for Predicting the Remaining Useful Life of Lithium-Ion Batteries

### v2

@@ 4.3 Validation of Prediction Model
Table 1: Performance comparison of deep learning models on the testing data of the second dataset.
Our Signal Preprocessing Model RMSE $\downarrow$ $\bm{R}^{2}$uparrow MAPE (%) $\downarrow$
W/o $\dot{Q}_{i}(Q,I)$ Attentional LSTM (A-LSTM) 174.83 0.84 8.14
CNN + A-LSTM 167.37 0.85 7.62
CNN + A-LSTM + ODE-A-LSTM 168.51 0.84 8.01
CNN + A-LSTM + ODE-LSTM (Ours) 163,31 0,85 7,54
With $\dot{Q}_{i}(Q,I)$ Attentional LSTM (A-LSTM) 168.09 0.87 7.47
CNN + A-LSTM 169.02 0.87 7.40
CNN + A-LSTM + ODE-A-LSTM 170.69 0.87 8.03
CNN + A-LSTM + ODE-LSTM (Ours) 158.09 0.89 6.95
@@
To assess the impact of various configurations on the prediction model performance, we conducted a series of experiments evaluating RMSE across different settings. These included activation functions (GeLU [16], HardSwish [2], ReLU [3], LeakyReLU), ODE integration schemes (RK4 [43], Midpoint [52], Heun2/3 [34], Euler [35]), convolutional kernel sizes, hidden layer sizes, and the selection of the relative time variable $R_{t}$ (RT), as defined in the A-LSTM paper [18]. Each configuration was isolated while keeping other components fixed to ensure fair comparison. The results, visualized in Figure 5, indicate that specific choices—such as using LeakyReLU as the activation function, Euler for ODE integration, a kernel size of $k=5$, a hidden size of 64, and first RT in A-LSTM block yielded lower RMSE values. These optimal settings were highlighted in the figure to emphasize their contribution to improved predictive accuracy. This analysis validates the importance of fine-tuning model configurations to achieve optimal performance in the target task.
@@
Table 2 presents a performance comparison between the proposed model and several baseline methods, including traditional machine learning models and the deep learning-based approach proposed in [24], across two battery datasets. On the first dataset, the proposed model significantly outperforms both ElasticNet [11] and XGBoostRegression [44], achieving the lowest RMSE of 101.59, the highest $R^{2}$ score of 0.82, and a substantially reduced MAPE of 8.29%. Similarly, on the second dataset, the proposed model surpasses all other methods, including the baseline model from [24], with an RMSE of 158.09, an $R^{2}$ score of 0.89, and the lowest MAPE of 6.95%. These results highlight the effectiveness and generalizability of the proposed approach across different data distributions.
Table 2: Comparison of the proposed approach with machine learning and existing approaches on both the first and second datasets.
@@
Dataset Method RMSE $\downarrow$ $\bm{R}^{2}$uparrow MAPE (%) $\downarrow$
First dataset ElasticNet 319.98 -0.20 26.70
XGBoostRegression 156.33 0.70 12.19
Our model 101.59 0.82 8.29
Second dataset ElasticNet 575.54 -0.20 26.67
XGBoostRegression 169.88 0.87 7.37
[24] 186 0.804 8.72
Our model 158.09 0.89 6.95
4.4 Extended Validation of Transfer Learning on the Prediction Model
@@
To evaluate the generalization capability of the proposed model across different battery datasets, we conducted a series of transfer learning experiments as detailed in Table 3. In cases involving upper and lower half splits, the training data were divided into two groups of battery cells based on cycle life: cells with a cycle life below the median of the cycle lives of the training set were assigned to the lower half, while those above the median were placed in the upper half. These experiments explore various scenarios in which the model is pre-trained on one dataset and fine-tuned on another, simulating real-world applications where labeled data in the target domain is scarce.
@@
Table 3: Configuration of transfer learning strategies for each case, showing the datasets used for pre-training and fine-tuning with their respective testing sets.
@@ 8 Entire training data of the first dataset Entire training data of the second dataset
First, we use Case 8 to investigate the impact of different block-freezing strategies during transfer learning, shown in Table 4. Notably, freezing only the A-LSTM block yields the best overall performance, suggesting that this component captures more transferable temporal features. In contrast, freezing all three blocks (CNN, A-LSTM, and ODE-LSTM) results in severe performance degradation, underscoring the importance of model flexibility during fine-tuning. Based on these findings, we adopt this block-freezing strategy of freezing only A-LSTM block for the next experiment, as shown in Table 5.
Table 4: Performance comparison under various block-freezing strategies applied durinng transfer learning in Case 8 (see Table 3).
Transfer Learning Strategy RMSE $\downarrow$ $\bm{R}^{2}$uparrow MAPE (%) $\downarrow$
@@
The results, summarized in Table 5, demonstrate that direct transfer (Cases 2 and 6) without fine-tuning leads to poor performance, with high RMSE and negative $R^{2}$, indicating a significant mismatch in data distributions between the two datasets. However, introducing fine-tuning with partial data from the target domain (Cases 3.1, 3.2, 7.1, and 7.2) noticeably improves model accuracy, with Case 7.2 achieving a strong balance (RMSE = 201.63, $R^{2}=0.81$, MAPE = 8.63%). Full fine-tuning on the target domain after pre-training (Cases 4 and 8) yields performance close to or even better than training directly on the target dataset (Cases 1 and 5), with Case 8 achieving the best results on the second dataset (RMSE = 157.92, $R^{2}=0.88$, MAPE = 6.88%).
Table 5: Performance comparison of our proposed model on different training strategies with Transfer Learning on both first and second dataset.
Case RMSE $\downarrow$ $\bm{R}^{2}$uparrow MAPE (%) $\downarrow$
@@
5 Conclusion
This study presents an robust approach for predicting the RUL of lithium-ion batteries by integrating an advanced signal processing pipeline with a novel deep learning prediction model. The signal preprocessing pipeline, comprising Capacity Interpolation and Denoising, Statistical Feature Extraction, Delta Feature Computation, and Feature Fusion, effectively captures the degradation dynamics over charge-discharge cycles. The proposed prediction model, combining CNN, A-LSTM, and ODE-LSTM blocks, demonstrates strong capabilities in modeling both discrete and continuous time behaviors of battery aging. Notably, the integration of our proposed CNN and ODE-LSTM block enhances the prediction model’s ability. The prediction model is further evaluated using transfer learning across different learning strategies and source data partitioning scenarios. Results indicate that the model maintains robust performance even when trained on limited data, underscoring its adaptability and effectiveness in data-constrained environments. Experimental validation on two large-scale datasets confirms that our method significantly outperforms a baseline deep learning approach and traditional machine learning techniques. These results highlight the potential of the proposed approach for practical deployment in predictive maintenance systems, contributing to improved reliability and longevity of battery-powered applications. Future work will explore the generalizability of the method across different battery chemistries and usage scenarios. Additionally, efforts will be directed toward real-time deployment on edge devices.

### v3

@@ IV-B Evaluation of Signal Preprocessing
TABLE I: Performance comparison of deep learning models on the testing data of the second dataset.
Our Signal Preprocessing Model RMSE $\downarrow$ $\bm{R}^{2}$ $\uparrow$ MAPE (%) $\downarrow$
W/o $\dot{Q}$ A-LSTM 174.83 0.84 8.14
A-LSTM + CNN 167.37 0.85 7.62
A-LSTM + CNN + ODE-A-LSTM 168.51 0.84 8.01
A-LSTM + CNN + LSTM 166.50 0.85 7.70
A-LSTM + CNN + ODE-LSTM (Ours) 163,31 0,85 7,54
With $\dot{Q}$ A-LSTM 168.09 0.87 7.47
A-LSTM + CNN 169.02 0.87 7.40
A-LSTM + CNN + ODE-A-LSTM 170.69 0.87 8.03
A-LSTM + CNN + LSTM 165.00 0.87 7.30
A-LSTM + CNN + ODE-LSTM (Ours) 158.09 0.89 6.95
Figure 4 (left) compares several reparameterizations: $\dot{Q}_{i}(I,Q)=\operatorname{Interp1D}(I_{i,j},Q_{i,j})\big|_{I_{k}}$, $\dot{V}_{i}(V,Q)=\operatorname{Interp1D}(V_{i,j},Q_{i,j})\big|_{V_{k}}$, $\dot{I}_{i}(I,Q)=\operatorname{Interp1D}(I_{i,j},Q_{i,j})\big|_{I_{k}}$, $\dot{Q}_{i}(Q,V)=\operatorname{Interp1D}(Q_{i,j},V_{i,j})\big|_{Q_{k}}$, $\dot{V}_{i}(V,I)=\operatorname{Interp1D}(V_{i,j},I_{i,j})\big|_{V_{k}}$, and $\dot{I}_{i}(I,V)=\operatorname{Interp1D}(I_{i,j},V_{i,j})\big|_{I_{k}}$. Our proposed feature is $\dot{Q}_{i}(I,Q)$. Note that all these derived features are applied during the signal preprocessing stage, up to and including the delta-feature computation. The results show that using only derived features yields the highest RMSE (252.52), indicating that relying solely on derived features leads to poor predictive performance. In contrast, the feature combination $I,V,Q,\dot{Q}(I,Q)]$ achieves the lowest RMSE (162.37), outperforming all other configurations. The strong contribution of $\dot{Q}$ underscores the effectiveness of Capacity Derivative Tracking for modeling discharge data: by aligning capacity values with corresponding current measurements, it provides a more consistent representation of the current–capacity relationship across cycles and improves predictive accuracy.
In the delta feature computation step, the difference in extracted features between each cycle and the one occurring delta steps earlier is computed. The choice of the delta parameter is important, as it determines the temporal resolution of feature variation. To assess its impact, we evaluated delta values ranging from 1 to 10, as shown in Figure 4 on the right. Among these, a delta value of 9 achieved the best performance, yielding the lowest RMSE of 162.37.
IV-C Validation of Prediction Model
As shown in Table I, the comparison between using only statistical features ($I$, $V$, $Q$) without $\dot{Q}$ and using combined fused features ($I$, $V$, $Q$, $\dot{Q}$) demonstrates that the proposed signal preprocessing pipeline—enhanced with the derived capacity feature $\dot{Q}$—consistently improves performance across all evaluated models. Notably, our prediction model, combining CNN, A-LSTM, and ODE-LSTM, achieves the best performance with the lowest RMSE of 158.09, highest $R^{2}$ score of 0.89, and lowest MAPE of 6.95%. These results confirm the effectiveness of integrating tailored CNN and ODE-LSTM blocks with A-LSTM, particularly when enriched by the derived capacity feature. Moreover, the results indicate that incorporating ODE into A-LSTM (ODE-A-LSTM) does not yield superior performance compared to the use of standard LSTM in this context.
@@ IV-C Validation of Prediction Model
To assess the impact of various configurations on the prediction model performance, we conducted a series of experiments evaluating RMSE across different settings. These included activation functions (GeLU [lee2023mathematical], HardSwish [avenash2019semantic], ReLU [banerjee2019empirical], LeakyReLU), ODE integration schemes (RK4 [tay2012spreadsheet], Midpoint [yu2019properties], Heun2/3 [saadoon2025numerical], Euler [salleh2012ordinary]), convolutional kernel sizes, hidden layer sizes, and the selection of the relative time variable $R_{t}$ (RT), as defined in the A-LSTM paper [liu2020anomaly]. Each configuration was isolated while keeping other components fixed to ensure fair comparison. The results, visualized in Figure 5, indicate that specific choices—such as using LeakyReLU as the activation function, Euler for ODE integration, a kernel size of $k=5$, a hidden size of 64, and first RT in A-LSTM block yielded lower RMSE values. The optimal settings consistently yield the lowest RMSE, indicating that these configurations better capture battery degradation signatures and, in turn, improve RUL prediction accuracy.
@@
Table II compares the proposed hybrid framework with several baselines: traditional machine learning models (ElasticNet [hans2011elastic], XGBoostRegression [wang2022xgboost]), Transformer-based deep learning model [vaswani2017attention], and current state-of-the-art (SOTA) approaches [ma2022real, ge2024structural], across two battery datasets. For the traditional machine learning models, the input data has been processed using our proposed signal preprocessing pipeline. On the first dataset, the proposed model significantly outperforms other models ElasticNet, XGBoostRegression, and the Transformer, achieving the lowest RMSE of 101.59, the highest $R^{2}$ score of 0.82, and a substantially reduced MAPE of 8.29%. Similarly, on the second dataset, the proposed model surpasses all other methods, including the state-of-the-art baselines from [ma2022real, ge2024structural], with an RMSE of 158.09, an $R^{2}$ score of 0.89, and the lowest MAPE of 6.95%. These results highlight the effectiveness and generalizability of the proposed approach across different data distributions. Compared with the two SOTA approaches [ma2022real, ge2024structural], which use a simple machine learning model and a shallow CNN that is insufficient to capture complex battery degradation dynamics, our prediction model adopts a hybrid architecture: a CNN + A-LSTM branch for characteristic feature extraction, an LSTM branch to capture temporal order, and an ODE-based component to model continuous degradation behavior. Although the resulting architecture is more expressive, the signal preprocessing pipeline reduces the small input size, keeping the overall model complexity compatible with practical hardware constraints (as detailed in Table III).
TABLE II: Performance comparison between the proposed approach, traditional machine learning models, and the state-of-the-art methods [ge2024structural, ma2022real] on the second dataset.
@@
Dataset Signal Preprocessing Prediction Model RMSE $\downarrow$ $\bm{R}^{2}$ $\uparrow$ MAPE (%) $\downarrow$
First dataset Ours ElasticNet 319.98 -0.20 26.70
Ours Transformers 193,83 0,47 16,59
Ours XGBoostRegression 156.33 0.70 12.19
Ours Ours 101.59 0.82 8.29
Second dataset Ours ElasticNet 575.54 -0.20 26.67
Ours Transformers 196,75 0,79 9,16
Ours XGBoostRegression 169.88 0.87 7.37
[ma2022real] [ma2022real] 186 0.804 8.72
[ge2024structural] [ge2024structural] 192.17 0.858 -
Ours Ours 158.09 0.89 6.95
Figure 7 compares the proposed ODE-LSTM with a standard LSTM. Unlike the standard LSTM, the ODE-LSTM introduces a continuous-time hidden state transition, enabling the latent dynamics to evolve smoothly between discrete observations. This results in more stable predictions across consecutive cycles, avoiding the abrupt fluctuations observed with the standard LSTM.
(a) With ODE-LSTM
(b) With LSTM
Figure 7: Comparison of our prediction model using ODE-LSTM and LSTM.
Figure 8 shows the RMSE distributions across cells on the testing data of the first and the second dataset. Bars indicate the empirical density, the solid curve is a Kernel Density Estimate (KDE) [chen2017tutorial] and the dashed curve is a fitted normal. The errors concentrate at low–to–moderate RMSE with a right-skewed tail driven by a few difficult cases. On the first dataset, the mean RMSE is 158.09 and most cells fall below RMSE=200, indicating accurate and stable RUL prediction with occasional outliers. On the second dataset, the mean RMSE decreases to 101.59, and the narrower histogram indicates improved consistency across cells.
(a) First Dataset
(b) Second Dataset
Figure 8: Across-cell RMSE distributions on the test sets
The results in Table III demonstrate that the proposed model is computationally efficient and suitable for real-world deployment. With only 1.53M parameters and a compact model size of 5.84 MB, the architecture has a small memory footprint that can be deployed on resource-constrained devices. Furthermore, on an AMD Ryzen 5 CPU, the measured inference latency for processing a single sample is below 4 ms (p90), with a throughput of nearly 287 inferences per second on a single core. This indicates that the model can deliver real-time predictions without requiring specialized hardware. Overall, these results validate that the proposed method achieves both high accuracy and practical efficiency, supporting its applicability for real-world RUL prediction tasks.
TABLE III: Computational efficiency of the proposed model.
Metric Value
Parameters 1,531,521
Model size (FP32) 5.84 MB
CPU latency (p50) 3.49 ms
CPU latency (p90) 4.12 ms
Mean latency 3.64 ms
Throughput 286.9 inferences/s
IV-D Extended Validation of the Prediction Model for Early Cycle Life Prediction
Figure 9: RMSE of cycle life early prediction (in cycles).
Figure 9 shows the RMSE of early cycle-life prediction on the first dataset. The bars represent the per-cell RMSE, the dashed line indicates our mean RMSE across all cells (181.11 cycles), and the dotted line marks the performance of the baseline model [severson2019data] (the original paper for the first dataset) at 214 cycles (as reported in Table 1 of [severson2019data]). Although we use the same experimental setup as [severson2019data], including the same data splitting, their model uses the first 100 cycles as input, whereas our model uses only the first 30 cycles. Despite this, our model still outperforms the baseline, demonstrating its strong overall performance for both RUL estimation and early cycle-life prediction.
IV-E Extended Validation of Transfer Learning on the Prediction Model
@@
To evaluate the generalization capability of the proposed model across different battery datasets, we conducted a series of transfer learning experiments as detailed in Table IV. In cases involving upper and lower half splits, the training data were divided into two groups of battery cells based on cycle life: cells with a cycle life below the median of the cycle lives of the training set were assigned to the lower half, while those above the median were placed in the upper half. These experiments explore various scenarios in which the model is pre-trained on one dataset and fine-tuned on another, simulating real-world applications where labeled data in the target domain is scarce.
@@
TABLE IV: Configuration of transfer learning strategies for each case, showing the datasets used for pre-training and fine-tuning with their respective testing sets.
@@ 8 Entire training data of the first dataset Entire training data of the second dataset
First, we use Case 8 to investigate the impact of different block-freezing strategies during transfer learning, shown in Table V. Notably, freezing only the A-LSTM block yields the best overall performance, suggesting that this component captures more transferable temporal features. In contrast, freezing all three blocks (CNN, A-LSTM, and ODE-LSTM) results in severe performance degradation, underscoring the importance of model flexibility during fine-tuning. Based on these findings, we adopt this block-freezing strategy of freezing only A-LSTM block for the next experiment, as shown in Table VI.
TABLE V: Performance comparison under various block-freezing strategies applied during transfer learning in Case 8 (see Table IV).
Transfer Learning Strategy RMSE $\downarrow$ $\bm{R}^{2}$ $\uparrow$ MAPE (%) $\downarrow$
@@
The results, summarized in Table VI, demonstrate that training on one dataset and evaluating on the other without transfer learning (Cases 2 and 6) yields markedly poor performance, with high RMSE and negative $R^{2}$, indicating a significant mismatch in data distributions between the two datasets.
However, introducing fine-tuning with partial data from the target domain (Cases 3.1, 3.2, 7.1, and 7.2) noticeably improves model accuracy, with Case 7.2 achieving a strong balance (RMSE = 201.63, $R^{2}=0.81$, MAPE = 8.63%). Full fine-tuning on the target domain after pre-training (Cases 4 and 8) yields performance close to or even better than training directly on the target dataset (Cases 1 and 5), with Case 8 achieving the best results on the second dataset (RMSE = 157.92, $R^{2}=0.88$, MAPE = 6.88%).
TABLE VI: Performance comparison of our proposed model on different training strategies with Transfer Learning on both first and second dataset.
Case RMSE $\downarrow$ $\bm{R}^{2}$ $\uparrow$ MAPE (%) $\downarrow$
@@
IV-F Discussion
In this study, we focused on evaluating the effectiveness of the proposed model under multiple charge and discharge profiles, while the batteries in the laboratory LFP/graphite lithium-ion datasets were tested under controlled, constant ambient temperature conditions. The results demonstrate that the model performs well across various charging conditions in the first dataset and across various discharging conditions in the second dataset. In future work, we plan to extend this study by (i) incorporating both cell and ambient temperature variations using real-world EV driving datasets, and (ii) testing additional battery datasets with different positive electrode materials, such as nickel–manganese–cobalt oxides (NMC), to further assess the chemical generalization capability of the model.
Furthermore, the current model utilizes only voltage and current signals as inputs. In subsequent research, we aim to integrate additional features derived from graph-based representations [zhao2025new, he2025dcaggcn]. For instance, in an EV battery pack, each sensor corresponding to a battery cell can be defined as a node, enabling the construction of a graph or hypergraph that captures inter-cell coupling effects (e.g., temperature correlations, shared cooling paths). These graph-structured features will then be combined with the proposed prediction model to enhance spatial–temporal degradation modeling.
Moreover, the proposed model has been evaluated using transfer learning between two datasets: the first containing various charge profiles and the second containing various discharge profiles. In future work, we will (i) extend this evaluation to cross-operating-condition transfer using real-world EV data (e.g., driving scenarios with acceleration, braking, and regenerative events), (ii) investigate the impact of different end-of-life thresholds (e.g., 70% vs. 80% capacity) on prediction accuracy, and (iii) explore uncertainty-aware or heteroscedastic loss functions to better characterize prediction confidence, particularly in late high-risk stages.
Finally, the predicted battery RUL provides actionable insights for optimizing charging and discharging strategies, scheduling battery replacement or recycling, and ultimately extending battery lifespan while reducing user anxiety in electric vehicle applications. In future work, we also plan to study how the proposed model can be embedded into on-board battery management systems with realistic computational and memory constraints.
V Conclusion
This study presents an robust approach for predicting the RUL of lithium-ion batteries by integrating an advanced signal processing pipeline with a novel prediction model. The signal preprocessing pipeline, comprising Capacity Interpolation and Denoising, Statistical Feature Extraction, Delta Feature Computation, and Feature Fusion, effectively captures the degradation dynamics over charge-discharge cycles. The proposed prediction model, combining CNN, A-LSTM, and ODE-LSTM blocks, demonstrates strong capabilities in modeling both discrete and continuous time behaviors of battery aging. Notably, the integration of our tailored CNN and ODE-LSTM block enhances the prediction model’s ability. The prediction model is further evaluated using transfer learning across different learning strategies and source data partitioning scenarios. Results indicate that the model maintains robust performance even when trained on limited data, underscoring its adaptability and effectiveness in data-constrained environments. Experimental validation on two large-scale datasets confirms that our method significantly outperforms baseline deep learning approaches and traditional machine learning techniques, indicating good adaptability in data-constrained settings. Future work will port the prediction model to ExecuTorch for on-device inference on edge hardware, enabling practical deployment in predictive maintenance systems. We will also evaluate the method’s generalizability across different battery chemistries and usage scenarios, and pursue real-time edge deployment.
Preprint Availability
A preprint of this manuscript is available at: https://arxiv.org/pdf/2505.16664